You are considering purchasing a 50,000 rentable sf office building. You project a

      $30.50 full service rental rate (operating expenses are projected to be $8.50 in the first

year). If you project 3% annual rent escalation and 4% annual operating expense escalation and it costs 13,000,000, what is the NPV after 5 years assuming an 8% annual discount rate and a sale at the end of 5 years using an 8% exit cap rate.

Sale expenses are 6% of sales price. Assume a 5% vacancy and bad debt factor and assume no debt – all cash transaction.
Answer #1
Year 0 1 2 3 4 5
Initial Investment -13000000
Annual Rent 0 50000*30.5*.95 = 1448750 1448750*1.03 = 1492212.5 1536979 1583088 1630580
Operating expense 0 50000*8.5 = 425000 425000*1.04 = 442000 459680 478067 497190
Sale 0 0 0 0 0 13000000*1.08^5 = 19101264
Net cash flow -13000000 1023750 1050212 1077299 1105021 20234654
Discount factor 1/1.08^0 = 1 1/1.08^1 = 0.926 0.857 0.794 0.735 0.681
Present value -13000000 947992.5 900031.7 855375.4 812190.4 13779799

NPV = Sum of present values = 4,295,389.374